OpenStudy (anonymous):

In a Triangle, I want to know the angle alpha: The adjacent side of alpha is 14, the opposite side is 1. How do I calculate alpha in degrees? Will give medal.

OpenStudy (mathstudent55):

The measures of the length of one side and one angle are not enough information to solve a triangle.

OpenStudy (anonymous):

maybe i expressed myself not clearly, sorry for that. I'll edit the question. Please help.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

This is what you wrote. There isn't enough information given to find anything else.

OpenStudy (mathstudent55):

Is this a right triangle?

OpenStudy (anonymous):

yes, I forgot that. Can it be solved then?

OpenStudy (anonymous):

figuered it out myself, answer is 4.086.

OpenStudy (mathstudent55):

 The function that relates the opposite leg to the adjacent leg is the tangent. \( \tan \alpha = \dfrac{opp}{adj} \) \( \tan \alpha = \dfrac{1}{14} \) \( \alpha = \tan^{-1} \dfrac{1}{14} \) \( \alpha = 4.086^o\) You are correct.